A cook puts 9.80 g of water in a 2.00-L pressure cooker that is then warmed to 550°C. What is the pressure inside the container? (Assume room temperature is 20°C.)
Answer in kPa
Answer #1

From ideal gas equation,

P2 = P1 T2/T1

P2 = 823 x 101325/293 = 2.85 x 10^5 Pa

Also,

PV = nRT

P = (9.80/18) x 8.314 x 823/0.002 = 1.86 x 10^6 Pa

Total pressure, P = 1.86 x 10^6 + 2.85 x 10^5 = 2.15 x 10^6 Pa = 2147.66 kPa

Comment in case any doubt please rate my answer....